OpenStudy (anonymous):

Which of the following choices is the common difference between the terms of this arithmetic sequence? 4x + 7y, 8x + 2y, 12x + -3y, 16x - 8y, 20x - 13y, ... A. 5x + 4y B. 4x - 5y C. -5x - 4y D. 5x - 4y E. 4x + 5y

OpenStudy (danjs):

how does each term change for every increase 4x to 8x is an increase of +4x

OpenStudy (danjs):

7y to 2y is a decrease of -5y

OpenStudy (danjs):

the next terms follow the same process

OpenStudy (danjs):

n2 - n1 = 8x + 2y 4x + 7y - --------------- 4x -5y

OpenStudy (danjs):

= n3-n2 = n4 - n3 = n5 - n4

OpenStudy (anonymous):

would it be C?

OpenStudy (danjs):

You subtract the nth term from the (n+1) term ... 8x+2y 4x+7y - ---------- 4x - 5y <----------common differ

OpenStudy (anonymous):

Then it would be B

OpenStudy (danjs):

Take any of the terms, then subtract from that the term before it.

OpenStudy (danjs):

right